# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

### v1

@@ 3.1.2 Bottom-up Engineering Modelling of RECs
Modeling these entities at a fine-grained level allows a detailed assessment of how different categories of users interact with their renewable energy systems over time. This insight is crucial for calculating quantities such as self-consumed energy and injected energy into the grid, which are useful for determining the change in energy volumes and thus for estimating economic market variables. An overview of the proposed bottom-up engineering methodology is shown in Fig. 2, which is explained in the following paragraphs.
@@
Consumption patterns for each prosumer were modeled using nPro (Wirtz, 2023), a load profiling tool capable of generating hourly demand curves based on user-specific behavior. To explore different energy scenarios, the load profiles were scaled to achieve three predefined self-consumption levels: 45%, 50%, and 55%.
This approach resulted in the generation of annual hourly-based time series (8,760 values) for each prosumer/producer category and each scenario, capturing PV generation, electricity consumption, self-consumed energy and energy injected into the grid. With seven market zones and three levels of self-consumption considered, the final output of the engineering model consists of twenty-one structured datasets. Each dataset includes the energy behavior of the four prosumer categories and the standalone PV systems, whose entire production is fed into the grid. These profiles are then used as input for the subsequent projection step, which is described in Section 3.1.3.
@@
To characterize the hourly production of PV systems across Italy, we made use of the PVGIS platform developed by the European Commission’s Joint Research Centre (PVGIS, 2022). For each of the seven electricity market zones we selected a reference city as defined above, and extracted location-specific PV yield data. Rather than relying on generic assumptions, we configured each PV system in PVGIS under optimal operating conditions. This meant that, for each site, the tilt angle of the PV modules was set to the value maximizing annual output, the azimuth orientation was chosen to face due south, which typically yields the best yearly performance in Italy and the PV technology considered was crystalline silicon, the most widespread module type. Also, system losses, such as reflecting factors, inverter efficiency, temperature effects, cable losses, dust, and shading, were included using an average value of 15% (Ogliari et al., 2023).
@@
  • $Y^{PV,one}_{p,z}(t)$ [kWh/kWp] is the hourly-based annual yield profile of the reference city from PVGIS of category p in electricity market zone z for one prosumer/producer
  • $E^{PV,one}_{p,z}(t)$ [kWh] is the hourly-based annual profile of energy produced by the PV system of category p in electricity market zone z for one prosumer/producer
@@
However, to account for the likely evolution of REC configurations in the coming years, we also explored additional scenarios featuring higher self-consumption levels. These extended cases are intended to emulate the effect of widespread adoption of energy storage systems, such as residential or commercial batteries. By shifting consumption toward daylight hours or enabling the deferred use of solar generation, energy storage reduces the amount of surplus energy injected into the grid, thus increasing the local use of renewable electricity. Although modelling of storage systems is not explicitly included in this framework, these scenarios of increased self-consumption provide an indication of their aggregate effect on energy flows. This allows an exploratory assessment of how the progressive penetration of storage technologies may alter the energy balance and the economic impact of RECs on the market.
@@
Once both the hourly PV generation and the electricity demand profiles were established for each prosumer/producer category and scenario, the next step involved calculating the two profiles that define the interaction between local generation and the grid: self-consumed energy and surplus energy injected into the main grid. These quantities were derived directly from the hourly time series of PV production $E^{PV,one}_{p,z}(t)$ and electrical load $E^{L,one}_{p,z}(t)$ of category p in electricity market zone z computed as explained in the previous section. In particular, the self-consumed energy $E^{self,one}_{p,z}(t)$ of category p in electricity market zone z for one prosumer/producer at each hour is computed as in (5):
@@
Figure 3: Hourly energy profiles for Milan (NORD) over a full year for four prosumer user categories: residential, school, commercial, and office. Each subplot shows PV generation (yellow line), electricity consumption (blue line), and energy injected into the grid (green line).
@@ 3.1.3 Scenarios and Projection of REC Deployment
Table 2: List of simulated scenario assumptions. sc = self consumption.
@@
10 Mixed scenario 1 sc_mix1.2027 2027 Public: 50% Residential: 45% SME: 55% NPO: 50% Standalone: no sc No
11 Mixed scenario 2 sc_mix2.2027 2027 Public: 55% Residential: 55% SME: 55% NPO: 55% Standalone: no sc Yes
Within each of these trajectories, scenarios are further differentiated based on assumed self-consumption rates, defined as the share of generated renewable energy consumed by prosumers within a REC rather than injected into the grid. This parameter is a critical driver in the modelling framework, as it reflects the efficiency and “virtuosity” of RECs in optimizing on-site energy use - thus affecting both demand-side and supply-side market dynamics. Additionally, two “mixed scenarios” are included to reflect more granular assumptions based on prosumer-category segmentation, as identified through our empirical mapping of REC profiles. These scenarios apply differentiated self-consumption rates by prosumer category (e.g., residential, public, SME, NPO), thereby enhancing the realism and policy relevance of the simulated outcomes. A summary of all scenarios is provided in Table 2.
The inputs for the scenario-based projections were provided by the engineering model that yielded two key variables related to individual prosumers/producers within a REC: annual energy self-consumed, $E_{self}(t)$, and the energy injected into the grid, $E_{export}(t)$, by typical prosumer categories and standalone PV systems. To represent the additional dimensionality of these variables, namely prosumer category and market zone as well as their individual nature, we denote them as $E_{p,z}^{\mathrm{self.REC}}(t)$ and $E_{p,z}^{\mathrm{export.ind}}(t)$ further in text. These values for self-consumed and injected energy were projected for both the ex-post (2024) and predicted (2027) deployment scenarios (see Table 2), using a set of parameters derived from the collected database.
The first projection parameter, “Zonal Shares of RECs,” represents the percentage distribution of RECs across electricity market zones:
@@
The total REC deployed PV capacity per prosumer category $p$ and zone $z$ was calculated for each scenario using the equation:
@@
The number of PV plants per prosumer category $p$ and market zone $z$ can be calculated using:
@@
  • $E_{p,z}^{\mathrm{self,one}}(t)$ is the energy self-consumed during hour $t$ by one prosumer of category $p$ in market zone $z$.
@@ 3.2 Second stage: Economic modeling
To assess the short-run economic impact of RECs on the Italian wholesale power market, we apply an empirical hour-by-hour counterfactual simulation similarly to the work by Beltrami et al. (2021), further extended by Beltrami (2024). This methodology builds on publicly available data from Gestore del Mercato Elettrico – GME – namely, “Offerte Pubbliche” (or public offers) – to reconstruct the merit-order demand and supply curves for each of the Italian electricity market zones and each hourly settlement period. From now onwards, in order to maintain the focus on the synthetic approach, we employ a simplified notation which is consistent with the formal used in Subsection 3.1. For instance, $\Delta Q_{\text{REC,d}}$ corresponds to $E_{z}^{self}(t)$, while $\Delta Q_{\text{REC,s}}$ corresponds to $E_{z}^{exp}(t)$ in Eq. (13) and (14). The reference to time and zone is omitted. The strategy holds for all zones and settlement periods. Our counterfactual simulation aims to model the hypothetical configuration of the day-ahead market in the absence of REC-driven generation and self-consumption.
This is a formal description of the theoretical approach we adopted. We consider an electricity market where, given a price level $Pr$, the demand curve is denoted by $D_{actual}(Pr)$ and the supply curve by $S_{actual}(Pr)$. These functions are derived from the aggregation of individual bids and offers submitted in a uniform-price day-ahead auction. The market equilibrium is determined by the intersection of these curves at the market-clearing price $Pr_{actual}$, such that:
$D(Pr_{actual})=S(Pr_{actual})=Q_{actual}$
where $Q_{actual}$ is the corresponding equilibrium quantity. In this setting, synthetic supply and demand curves can be constructed to represent a competitive benchmark, using observed bidding behavior or reconstructed marginal costs across all operators. These curves serve as the baseline for analyzing the effect of RECs.
@@
As stated above, RECs typically self-consume a share of their electricity production. This self-consumed electricity is not visible in the observed market demand. In a counterfactual scenario where RECs are absent, this hidden demand would need to be satisfied by the wholesale market.
Let $\Delta Q_{\text{REC,d}}$ denote the total quantity of REC self-consumption. To account for this, the demand curve is adjusted by shifting it horizontally to the right:
$D_{synt}(Pr)=D(Pr)+\Delta Q_{\text{REC,d}}$
@@
On the supply side, RECs may inject electricity into the market through surplus generation. Removing RECs implies that this contribution is also removed from the market supply. Let $\Delta Q_{\text{REC,s}}$ represent the total quantity of REC-generated electricity that would have been offered to the market. The adjusted supply curve is thus defined as:
$S_{synt}(Pr)=S(Pr)-\Delta Q_{\text{REC,s}}$
@@
$D_{synt}(Pr_{synt})=S_{synt}(Pr_{synt})$
@@
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}=S(Pr_{synt})-\Delta Q_{\text{REC,s}}$
@@
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}+\Delta Q_{\text{REC,s}}=S(Pr_{synt})$
@@
This dual (contemporaneous) adjustment is performed hourly and for each market zone, generating synthetic demand and supply curves representing a no-REC baseline condition. The resulting counterfactual price and quantity outcomes are then compared with the observed market outcomes to estimate the impact of REC operations.
We acknowledge that this simulation is based on several assumptions. First, the counterfactual is computed under a ceteris paribus condition, assuming all other market dynamics unchanged. [^22] This assumption is particularly relevant when comparing the effects of RECs on demand and supply across the different scenarios, and becomes evident mostly in interpreting the results from the Policy Scenario compared to BU and HW. In our setting, we implicitly assume that the installed capacity of other RES technologies (beyond RECs) remains fixed. This does not reflect the actual dynamics of the Italian power system, where RES capacity additions are currently progressing at sustained rates as well as does not reflect the demand growth. Nevertheless, this choice is consistent with our primary goal, i.e. to isolate the specific marginal contribution of REC deployment, without counfouding their individual effect with broader renewable expansion trends. Second, each market zone is treated as a “closed system”, abstracting from inter-zonal electricity flows. While this assumption may limit economic and policy implications, it remains defensible for two main reasons: (1) RES-generated power is typically prioritized in dispatch due to the merit-order principle, making its theoretical removal analytically legitimate; (2) the current scale of REC operations in Italy is still limited, rendering its impact negligible in terms of strategic bidding by large market players deeply affecting the market clearing dynamics. Overall, this synthetic control approach allows us to isolate and quantify the localized, short-term effects of REC deployment on market outcomes such as zonal clearing prices and traded quantities, thus offering a robust empirical basis for evaluating REC-driven welfare gains.

### v2

@@ 3.1.2 Bottom-up Engineering Modelling of RECs
Modeling RECs at a fine-grained level allows a detailed assessment of how different categories of users interact with their renewable energy systems over time. This insight is crucial for calculating quantities such as self-consumed energy and injected energy into the grid, which are useful for determining the change in energy volumes and thus for estimating economic market variables. An overview of the proposed bottom-up engineering methodology is shown in Fig. 2, which is explained in the following paragraphs.
@@
Consumption patterns for each prosumer were modeled using nPro (Wirtz, 2023), a load profiling tool capable of generating hourly demand curves based on user-specific behavior. To explore different energy scenarios, the load profiles were scaled to achieve three predefined self-consumption levels: 45%, 50%, and 55%. This approach resulted in the generation of annual hourly-based time series (8,760 values) for each prosumer/producer category and each scenario, capturing PV generation, electricity consumption, self-consumed energy and energy injected into the grid. With seven market zones and three levels of self-consumption considered, the final output of the engineering model consists of twenty-one structured datasets. Each dataset includes the energy behavior of the four prosumer categories and the standalone PV systems, whose entire production is fed into the grid. These profiles are then used as input for the subsequent projection step, which is described in Section 3.1.3.
@@
To characterize the hourly production of PV systems across Italy, we made use of the PVGIS platform developed by the European Commission’s Joint Research Centre (PVGIS, 2022). For each of the seven reference city, we extracted location-specific PV yield data. Rather than relying on generic assumptions, we configured each PV system in PVGIS under optimal operating conditions. This meant that, for each site, the tilt angle of the PV modules was set to the value maximizing annual output, the azimuth orientation was chosen to face due south, which typically yields the best yearly performance in Italy and the PV technology considered was crystalline silicon, the most widespread module type. Also, system losses, such as reflecting factors, inverter efficiency, temperature effects, cable losses, dust, and shading, were included using an average value of 15% (Ogliari et al., 2023).
@@
  • $E^{PV,one}_{p,z}(t)$ [kWh] is the hourly-based annual profile of energy produced by the PV system of category p in electricity market zone z for one prosumer/producer.
  • $Y^{PV,one}_{p,z}(t)$ [kWh/kWp] is the hourly-based annual yield profile of the reference city from PVGIS of category p in electricity market zone z for one prosumer/producer.
@@
However, to account for the likely evolution of REC configurations in the coming years, we also explored additional scenario featuring higher self-consumption levels (mixed scenario in Table 2). This extended case is intended to emulate the effect of widespread adoption of energy storage systems, such as residential or commercial batteries. By shifting consumption toward daylight hours or enabling the deferred use of solar generation, energy storage reduces the amount of surplus energy injected into the grid, thus increasing the local use of renewable electricity. Although modelling of storage systems is not explicitly included in this framework, this scenario of increased self-consumption provide an indication of their aggregate effect on energy flows. This allows an exploratory assessment of how the progressive penetration of storage technologies may alter the energy balance and the economic impact of RECs on the market.
@@
Once both the hourly PV generation and the electricity demand profiles were established for each prosumer/producer category and scenario, the next step involved calculating the two profiles that define the interaction between local generation and the grid: self-consumed energy and surplus energy injected into the main grid. These quantities were derived directly from the hourly time series of PV production $E^{PV,one}_{p,z}(t)$ and electrical load $E^{L,one}_{p,z}(t)$ of category p in electricity market zone z computed as explained in the previous sub-section. In particular, the self-consumed energy $E^{self,one}_{p,z}(t)$ of category p in electricity market zone z for one prosumer/producer at each hour is computed as in Eq.(4):
@@
Figure 3: Hourly energy profiles for Milan (NORD) over a full year for four prosumer user categories. Each subplot shows PV generation (yellow line), electricity consumption (blue line), and energy injected into the grid (green line).
@@
Figures 4 and 5 display the same hourly profiles albeit over an average day in January and April. We see that in January (Figure 4) the greatest self-consumption (intersection between yellow and blue bars) is exhibited by a school. Moreover, its load (blue bars) surpasses its production during midday off-peak hours (10 a.m.–15 p.m.), thereby potentially offsetting an extra generation from REC’s residential and commercial prosumers. The load profile of an office is very similar to the school’s profile, though on a smaller scale (vertical axes). In April, we see that the absence of heating and cooling needs entails a more moderate demand for energy. Consequently, a lot of excessive energy goes into the grid. In both graphs, we observe that residential prosumers exhibit much smaller energy flows and a different load profile than other prosumers. In addition to evening peak difference, a peak midday load of a school, commercial and office buildings occurs at 9 a.m.–12 a.m., whereas of a residential building at 13 p.m.–14 p.m., although more pronounced in January than in April.
Figure 4: Hourly energy profiles for Milan (NORD) over an average day in January for four prosumer user categories. Each subplot shows PV generation (yellow bars), electricity consumption (blue bars), and energy injected into the grid (green bars).
Figure 5: Hourly energy profiles for Milan (NORD) over an average day in April for four prosumer user categories. Each subplot shows PV generation (yellow bars), electricity consumption (blue bars), and energy injected into the grid (green bars).
@@ 3.1.3 Scenarios and Projection of REC Deployment
Table 2: List of simulated scenarios
@@
10 Mixed scenario sc_mix1.2027 2027 Public: 50% Residential: 45% SME: 55% NPO: 50% Standalone: no sc No
Within each of these trajectories, scenarios are further differentiated based on assumed self-consumption rates, defined as the share of generated renewable energy consumed by prosumers within a REC rather than injected into the grid. This parameter is a critical driver in the modelling framework, as it reflects the efficiency and “virtuosity” of RECs in optimizing on-site energy use - thus affecting both demand-side and supply-side market dynamics. Additionally, a “mixed scenario” is included to reflect more granular assumptions based on prosumer-category segmentation, as identified through our empirical mapping of REC profiles. These scenario applies differentiated self-consumption rates by prosumer category, thereby enhancing the realism and policy relevance of the simulated outcomes. A summary of all scenarios is provided in Table 2.
The scenario-based projection inputs were drawn from (i) the engineering model, which provided two key variables for one typical prosumer/producer within a REC—annual energy self-consumed and energy injected into the grid—and (ii) projection parameters derived from our database. The first projection parameter, “Zonal Shares of RECs,” represents the percentage distribution of RECs across electricity market zones:
@@
The third parameter, ”total REC deployed PV capacity per prosumer category $p$ and zone zz” was calculated for each scenario using the equation:
@@
The fourth parameter, the ”number of PV plants per prosumer category $p$ in market zone zz” can be calculated using:
@@
  • $Plant_{p,z}$} is the number of PV plants per prosumer category $p$ in market zone $z$.
@@
  • $E_{p,z}^{\mathrm{self,one}}(t)$ is the energy self-consumed during hour $t$ by one prosumer of category $p$ in market zone $z$,
  • $t$ is a specific hour in a year, $t\in T$ and $T=\{1,2,\ldots,8760\}$.
@@ 3.2 Second stage: Economic modeling
To assess the short-run economic impact of RECs on the Italian wholesale power market, we apply the empirical hourly counterfactual simulation proposed by Beltrami et al. (2021), further extended by Beltrami (2024). This methodology builds on publicly available data from Gestore del Mercato Elettrico – GME – namely, the “Offerte Pubbliche” – which report all price–quantity bids submitted to the Italian day-ahead auction. These data allow us to reconstruct the hourly merit-order demand and supply curves for each electricity market zone.
For a given hour and zone, each supply bid is characterized by an offered quantity $q_{i}^{s}$ and an associated bid price $p_{i}^{s}$, while demand bids are defined by requested quantities $q_{j}^{d}$ at bid prices $p_{j}^{d}$. Following the merit-order principle, supply bids are ranked in ascending price order and aggregated into a stepwise supply curve:
$S_{actual}(Pr)=\sum_{i:p_{i}^{s}\leq Pr}q_{i}^{s}$ (14)
Analogously, demand bids are ordered in descending price order and aggregated into a stepwise demand curve:
$D_{actual}(Pr)=\sum_{j:p_{j}^{d}\geq Pr}q_{j}^{d}$ (15)
By construction, both curves are piecewise-constant functions, reflecting the discrete nature of bidding in day-ahead auctions. The market-clearing price $Pr_{actual}$ is defined as the minimum price such that:
$S_{actual}(Pr_{actual})\geq D_{actual}(Pr_{actual}),$ (16)
with the corresponding cleared quantity given by $Q_{actual}=D_{actual}(Pr_{actual})$.
From now onwards, to maintain the focus on the synthetic counterfactual approach, we employ a simplified notation consistent with Subsection 3.1 In particular, $\Delta Q_{\text{REC,d}}$ corresponds to $E_{z}^{self}(t)$, while $\Delta Q_{\text{REC,s}}$ corresponds to $E_{z}^{exp}(t)$ in Eq. (12) and (13). [^23] Time and zonal indices are omitted for expositional clarity, while the strategy applies to all zones and settlement periods. Crucially, the counterfactual simulation aims to reproduce the hypothetical configuration of the day-ahead market in the absence of REC-driven generation and self-consumption. In this framework, the observed demand and supply curves, $D_{actual}(Pr)$ and $S_{actual}(Pr)$, constitute the empirical baseline from which synthetic curves are constructed to evaluate the equilibrium effects induced by RECs.
@@
As stated above, RECs typically self-consume a share of their electricity production. This self-consumed electricity is not visible in the observed market demand. In a counterfactual scenario where RECs are absent, this hidden demand would need to be satisfied by the wholesale market. Let then $\Delta Q_{\text{REC,d}}$ denote the total quantity of REC self-consumption. To account for this, the demand curve is adjusted by shifting it horizontally to the right:
$D_{synt}(Pr)=D(Pr)+\Delta Q_{\text{REC,d}}$ (17)
@@
On the supply side, RECs inject electricity into the market through surplus generation [^24] The GSE is responsible for informing the market operator, GME, about distributed generation volumes. The REC-driven quantities are then placed on the supply curve at zero price, along with the other distributed renewable generation.. Removing RECs implies that this contribution is also removed from the market supply. Let $\Delta Q_{\text{REC,s}}$ represent the total quantity of REC-generated electricity that would have been offered to the market. The adjusted supply curve is thus defined as:
$S_{synt}(Pr)=S(Pr)-\Delta Q_{\text{REC,s}}$ (18)
@@
$D_{synt}(Pr_{synt})=S_{synt}(Pr_{synt})$ (19)
@@
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}=S(Pr_{synt})-\Delta Q_{\text{REC,s}}$ (20)
@@
$D(Pr_{synt})+\Delta Q_{\text{REC,d}}+\Delta Q_{\text{REC,s}}=S(Pr_{synt})$ (21)
@@
This dual (contemporaneous) adjustment is performed hourly and for each market zone, generating synthetic demand and supply curves representing a no-REC baseline condition. The resulting counterfactual quantity outcomes are then compared with the observed market outcomes to estimate the impact of REC operations. To quantify the magnitude of REC effects on market outcomes, we define the hourly percentage impact on equilibrium quantities as the relative difference between the observed equilibrium quantity in the presence of RECs and the counterfactual equilibrium quantity in their absence. Formally, for each hour and market zone, the indicator is computed as:
$\text{Impact}_{Q,h}=\frac{Q_{actual,h}-Q_{synt,h}}{Q_{synt,h}}\times 100,$ (22)
where $Q_{actual,h}$ denotes the cleared quantity observed in the day-ahead market with RECs, while $Q_{synt,h}$ represents the synthetic equilibrium quantity obtained under the no-REC counterfactual scenario.
A positive value of the computed impact in Eq. (22) indicates that REC deployment is associated with higher market-cleared volumes relative to the counterfactual baseline, whereas a negative value signals a reduction in traded quantities, typically driven by higher levels of REC self-consumption. This metric is computed on an hourly basis and then analyzed across months, seasons, and policy scenarios, as reported in Section 4.
We acknowledge that this simulation is based on several assumptions. First, the counterfactual is computed under a ceteris paribus condition, assuming all other market dynamics unchanged. [^25] This assumption is particularly relevant when comparing the effects of RECs on demand and supply across the different scenarios, and becomes evident mostly in interpreting the results from the Policy Scenario compared to BU and HW. In our setting, we implicitly assume that the installed capacity of other RES technologies (beyond RECs) remains fixed. This does not reflect the actual dynamics of the Italian power system, where RES capacity additions are currently progressing at sustained rates as well as does not reflect the demand growth. Nevertheless, this choice is consistent with our primary goal, i.e. to isolate the specific marginal contribution of REC deployment, without counfouding their individual effect with broader renewable expansion trends. Second, each market zone is treated as a “closed system”, abstracting from inter-zonal electricity flows. While this assumption may limit economic and policy implications, it remains defensible for two main reasons: (1) RES-generated power is typically prioritized in dispatch due to the merit-order principle, making its theoretical removal analytically legitimate; (2) the current scale of REC operations in Italy is still limited, rendering its impact negligible in terms of strategic bidding by large market players deeply affecting the market clearing dynamics. Overall, this synthetic control approach allows us to isolate and quantify the localized, short-term effects of REC deployment on market outcomes such as zonal traded quantities and clearing prices, thus offering a robust empirical basis for evaluating REC-driven welfare gains.